he physician orders Gentamicin 5 mg/kg/dose. The patient weighs 165 pounds. Gentamicin is available as 20 mg/2 mL. How many mL will each dose contain?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug per dose.

       Formula = Desired dose/Dose in hand * Quantity.

   Here

          Ordered dose = 5 mg/kg/dose.

          Weight of the patient = 165 pounds.

       Convert pound to kg for calculation

                   1 pound = 0.454 kg.

        So

           Weight of the patient = 165 * 0.454

                                              = 74.9 kg.

Desired dose(D) = 5 * 74.9

                                         = 374.5 mg.

             Dose in hand (H) = 20 mg.

                     Quantity (Q) = 2 ml.

    As per above formula

          Volume(ml) of drug per dose = D/H * Q

                                                          = 374.5 / 20 * 2

                                                           = 18.73 * 2

                                                           = 37.5 ml per dose.